OpenStudy (anonymous):

Find the missing angle measure in QRS. Q = _____°, R = 60°, S = 20° A. 80 B. 100 C. 120 D. 60

OpenStudy (anonymous):

R+S+Q=180 60+20+Q=180 Q=180-80=100

OpenStudy (anonymous):

sum of three angles in a triangle is 180 degrees

OpenStudy (whpalmer4):

The interior angles of a polygon with n sides add up to (n-2)*180 (degrees). This is a triangle, 3 sides, n = 3, so the sum of the interior angles is (3-2)*180 degrees. You've already got 60+20 = 80 degrees accounted for, so the remaining angle is 180 degrees - 80 degrees = 100 degrees

OpenStudy (whpalmer4):

With a rectangle, there are 4 sides, so n = 4, and the interior angles add up to 2*180 degrees = 360 degrees.

OpenStudy (anonymous):

thanks!